If the in 2017 equals 200 and prices are 3% higher than in 2016what is the 2016 value of the CPI?
Answer #1

CPI in 2016 = CPI in 2017/(1+difference in price level)

= 200/(1+0.03) = 200/1.03 = 194.17 is the answer to this question